Rani ate 2/7 part of a cake while her brother Ravi ate 4/5 of the remaining. Part of the cake left is                                         

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Rani ate 2/7 part of a cake while her brother Ravi ate 4/5 of the remaining. Part of the cake left is 1/7.
Now, let us assume total part of cake be 1.
Then, from the question, given that Rani ate 2/7 part of a cake = 1 - (2/7)[
So, Ravi ate 4/5 of the remaining cake = (4/5) × (5/7)
= 4/7

Therefore, the part of the cake left = (5/7)–(4/7)
= 1/7

Thanks